# MC-NeRF: Multi-Camera Neural Radiance Fields for Multi-Camera Image Acquisition Systems

### v2

@@ 3.4.2   Coupling in camera parameters
$\tilde{t}_{3}$ $=1-t_{3}$
In Eq.(14) and Eq.(15), ${r_{ij}}$ represents element of the rotation matrix $R$ in 3D space, and $t_{i}$ represents the element of the translation matrix $T$ in 3D space. From Eq.(15), we can observe that $u_{0}$, $v_{0}$ and $t_{1}$, $t_{2}$, $t_{3}$ are strongly coupled. While we can obtain result for $\tilde{T}-{R^{-1}}T$ through optimization, distinguishing between $u_{0}$, $v_{0}$ and $T$ remains challenging.
A straightforward solution to address the coupling problem in joint optimization is by imposing additional constraints on the coupled variables. We introduce a constraint for the camera intrinsic parameters to separate intrinsic and extrinsic parameters from each other. In the collected calibration data, $Pack2$ provides comprehensive constraints on the camera intrinsic parameters, while $Pack1$ only constrains the world coordinate system with extrinsic parameters (the images in $Pack1$ cannot guarantee at least two Apriltags captured in each image). Moreover, the extrinsic parameters $RT_{local}^{i}$ generated by $Pack2$ are independent of the multi-camera acquisition system’s extrinsic parameters $RT_{global}^{i}$, which eliminates the impact of $RT_{global}^{i}$ while constraining the intrinsic parameters. The described procedure is shown in Fig.6.
In summary, as mentioned in Section 3.1, we require the intrinsic reprojection loss as a constraint to achieve the independent regression of coupled variables $u_{0}$, $v_{0}$ and $t_{1}$, $t_{2}$, $t_{3}$.
3.4.3   Neural Image Alignment (2D)
@@
As mentioned in BARF and L2G-NeRF, 2D neural image alignment problem can be described as: let $x$ be the 2D pixel coordinates, we aim to optimize a 2D neural field parameterized as the weights of a multilayer perceptron (MLP) ${f_{mlp}}$:
@@
while also solving for geometric transformation parameters as $T=[R{\rm{|}}T]$, and $[R{\rm{|}}T]$ denotes the rigid rotation and translation in 2D space.
@@
Based on Eq.(17) and Eq.(18), the regression for the intrinsic parameter $K$ can be equivalently represented as the regression for the linear transformation $L$. It’s important to note that in the experiments conducted by BARF and L2G, the term $T$ in Eq.(16) is defined as a homography matrix. However, since the transformation requirements for 3D space in the NeRF series methods involve only rotation and translation, we redefine $T$ as $[R{\rm{|}}T]$ instead. This adjustment reduces the number of regression parameters for each cropped patch from 8 ($T\in SL(3)$) to 3 ($T\in SE(2)$). Nonetheless, the inclusion of the linear transformation introduces an additional 4 parameters, resulting in a total of 7 regression parameters.
Figure 7: Scene styles and dataset structure. The left side shows the camera distribution for the different multi-camera systems and the eight objects of interest in our dataset. The right side illustrates the composition of our dataset. Within each style, there are eights objects correspond to five sets of data, including two sets of calibration ($Pack1$ and $Pack2$), as well as the training, validation, and test sets.
3.4.4   NeRF with Bundle Adjustment
Lin et al. [38] first proposed Bundle-Adjusting Neural Radiance Fields (BARF), which addresses the problem of achieving NeRF training without precise camera extrinsic parameters. We believe one of the key advantages of BARF is its independence from requiring extra information. It only needs simple adjustments to the coarse-to-fine NeRF architecture to achieve the joint optimization for camera pose and NeRF. The authors validated the proposed progressive alignment method on the image alignment task in 2D space, and extended its application to the 3D NeRF space.
We attempted to extend this method again by incorporating the regression of camera intrinsic parameters into the BARF framework. The training sequence map of our method is illustrated in Fig.5. During the camera parameters initialization stage, coarse camera parameters are obtained using the calibration data $Pack1$ and $Pack2$. In the global optimization stage, we employ the intrinsic reprojection loss branch to impose constraints on the principal point coordinates of the camera. This strategy effectively addresses the parameter coupling problem outlined in Section 3.4.2. Simultaneously, we leverage the progressive alignment method from BARF to optimize the extrinsic parameters of the multi-camera acquisition system. In the fine-tuning stage, we freeze the extrinsic parameters and only refine the intrinsic parameters to obtain the final results.
4   EXPERIMENTS
Figure 8: Comparison of rendering performance in NeRF with mixed intrinsic dataset and single intrinsic dataset.
The goal of the experiments is to answer the following questions:
  1. 1.
    Does NeRF still work effectively when each image corresponds to different intrinsic parameters? In comparison to the globally unique camera assumption in the original NeRF, what challenges do the multiple intrinsic parameters present to the NeRF series methods?
  2. 2.
    Does the Intrinsic Reprojection Loss branch proposed in Section 3.2 successfully yield accurate camera intrinsic parameters? Section 3.2 highlights that if the calibration points lie on the same plane, a degenerate case could occur, resulting in inaccurate estimation. Does this challenge still remain within the proposed network architecture?
  3. 3.
    Is the extrinsic regression computation introduced in Section 3.3 necessary? How does the performance of models such as BARF and L2G-NeRF, which can directly regress extrinsic parameters with accurate intrinsic parameters provided?
  4. 4.
    Compared to previous works, how does the proposed method perform in the 2D neural image alignment problem?
  5. 5.
    Does the approach introduced in Section 3.4 achieve global optimization of the network? Having obtained the intrinsic and extrinsic parameters of different cameras as initial values in Sections 3.2 and 3.3, how does the performance compare when directly optimizing NeRF using these initial parameters, without implementing global optimization?
4.1   Dataset and Hardware Platfrom
We designed four datasets based on commonly used multi-camera image acquisition systems. The camera distribution in these datasets includes halfball style (as employed by the original NeRF), ball style, room style, and array style. The datasets contain a total of 32 groups, with 8 different objects for each dataset style. The different sampling styles involve varying numbers of cameras, leading to different numbers of training images. For each style, the number of validation images is equal to the number of training images. The details of the datasets are illustrated on the right side of Fig.7.
It’s worth noting that due to the images being collected from cameras with different intrinsic parameters, determining which camera to use for generating the test data is challenging. To showcase the performance of the proposed method across various camera intrinsic parameters, we generate the 200 test images in each group by continuously varying the intrinsic parameters along the camera’s motion trajectory. Detailed sampling procedures and parameter variations have been provided on the project website.
The datasets are generated using Blender 3.3.3, and the intrinsic parameters for each camera within each style are different. For example, in the HalfBall Style acquisition system, there are a total of 100 cameras with varying intrinsic parameters. The training data is collected using these cameras according to their depicted positions. Similarly, the validation data is also collected using these cameras, but each camera’s position differs from that of the training set. The resolution of each image is $800\times 800$. All of our experiments were conducted on an NVIDIA RTX 3090 with 24GB, and PyTorch version is 1.12.1.
TABLE 2: Quantitative results of Exp.1.
$40^{\circ}$ $50^{\circ}$ $60^{\circ}$ $70^{\circ}$ $80^{\circ}$
@@ 4.1   Dataset and Hardware Platfrom
Mixed FOV 24.169 27.333 29.949 32.138 33.786
Distance 1.271 0.539 0.053 0.797 0.997
@@
Mixed FOV 0.819 0.899 0.939 0.963 0.975
Distance 0.039 0.014 0.003 0.006 0.005
[…]
Mixed FOV 0.147 0.069 0.037 0.026 0.016
Distance 0.039 0.013 0.003 0.009 0.002
4.2   Exp.1 Unique-Camera VS Multi-Cameras
In this experiment, our goal is to explore whether NeRF has the capability to utilize multiple intrinsic parameters simultaneously, which represents a foundational prerequisite for our method. We select cameras with FOVs of 40, 50, 60, 70, and 80 degrees to generate a mixed dataset. NeRF is then trained on this dataset with provided ground-truth intrinsic and extrinsic parameters. Simultaneously, we also conduct training of NeRF with each FOV data independently and compare the rendering performance with NeRF trained on the mixed dataset. Hence, this experiment contains a total of 6 datasets: 1 mixed dataset and 5 independent datasets. Each dataset has 84 images, and the extrinsic parameters are the same totally across all datasets.
The results are presented in Fig.8 where the top row shows ground-truth images captured at varying FOVs, indicated below each image. The second row displays results produced by the NeRF model trained on each fixed FOV dataset, and the third row exhibits results from the NeRF model trained on a mixed dataset. The numbers in the top-right corner of each result represent the number of images in the dataset with the current FOV. In the mixed dataset, there are 22, 12, 10, 24, and 16 images for FOVs of 40, 50, 60, 70, and 80 degrees, respectively.
Figure 9: Comparison results of regressing intrinsic parameters with different number of Apriltag in calibration data. With at least two Apriltags included in the calibration data, we can achieve the expected results. However, when including only one Apriltag, degenerate case is showed up, leading to poor results. Specifically, the black curve in $E1$ fails to converge to zero, while the blue and red curves both converge near zero.
The comparison results can be observed in Table 2. The $Distance$ item indicates the absolute difference between their respective indicators. As evident from both Table 2 and Fig.8, the models trained with different datasets exhibit nearly equivalent performance (in each column of both Fig.8 and Table 2). This observation has shown the capability of NeRF that can effectively process multi-camera images, when accurate camera intrinsic and extrinsic parameters are provided.
4.3   Exp.2 Intrinsic Parameters Regression
TABLE 3: Quantitative results of Exp.2.

### v4

@@ 3.2.1   Coupling in camera parameters
$\tilde{t}_{3}$ $=-t_{3}$
From the last line of Eq.(5), we can observe that the 2D-to-3D projection can be equated to operations based on $\tilde{R}$ and $(\tilde{T}-R^{-1}T)$. From Eq.(7), $\tilde{R}$ can be decomposed into a combination of scaling and rotation, which are decoupled.
However, for $(\tilde{T}-R^{-1}T)$, this item represents a translation operation, which is influenced by $R$, $K$ and $T$. Taking the first row of Eq.(8) as an example, we have:
$\tilde{t}_{1}r_{11}+\tilde{t}_{2}r_{12}+\tilde{t}_{2}r_{13}$ (9)
$\Rightarrow\alpha{u_{0}}+\beta$ ${v_{0}}+{r_{11}}{t_{1}}+{r_{12}}{t_{2}}+{r_{13}}{t_{3}}$
[…]
Since rotation and scaling are decoupled, we assume that we can obtain the correct optimization results for $R$ and $f_{x}$, $f_{y}$. This means that Eq.(9) contains five unknown variables: $u_{0}$, $v_{0}$ and $t_{1}$, $t_{2}$, $t_{3}$. According to Eq.(8), there are three constraint equations for the five parameters mentioned above, which indicates that it is not possible to accurately solve for each parameter. While we can obtain result for $(\tilde{T}-{R^{-1}}T)$ through optimization, distinguishing between $u_{0}$, $v_{0}$ and $T$ remains challenging.
In summary, $u_{0}$, $v_{0}$, and $T$ are coupled during the joint optimization of both intrinsics and extrinsics, making it impossible to obtain accurate values for each. As a result, we can achieve good rendering performance from the view of the training images, but rendering errors occur when generating new viewpoints.
3.2.2   Coupling in multi-camera systems
It is important to note that in Sec 3.2.1, we considered only a case of single image and a single camera. Next, we analyze the case of multiple images and multiple cameras. The existing methods can be categorized into the following three cases:
1) With intrinsics provided, joint optimization of extrinsics and NeRF, such as BARF, L2G-NeRF.
2) With global unique intrinsics, joint optimization of intrinsics, extrinsics and NeRF, such as SC-NeRF, NeRF--- -.
3) Joint optimization of intrinsics, extrinsics and NeRF, our goal and method.
Based on Eq.(8) and Eq.(9), $(\tilde{T}-R^{-1}T)$ have the following expression:
$\Rightarrow$ $\begin{bmatrix}{\alpha_{1}}{u_{0}}+{\beta_{1}}{v_{0}}+{r_{11}}{t_{1}}+{r_{12}}{t_{2}}+{r_{13}}{t_{3}}\\ {\alpha_{2}}{u_{0}}+{\beta_{2}}{v_{0}}+{r_{21}}{t_{1}}+{r_{22}}{t_{2}}+{r_{23}}{t_{3}}\\ {\alpha_{3}}{u_{0}}+{\beta_{3}}{v_{0}}+{r_{31}}{t_{1}}+{r_{32}}{t_{2}}+{r_{33}}{t_{3}}\\ \end{bmatrix}$ (10)
[…]
For the first case, when the intrinsics are provided, the number of unknown variables is reduced from five to three (only $t_{1}$, $t_{2}$, $t_{3}$), matching the number of equation constraints in Eq.(10), allowing for joint solving. This indicates that the extrinsics can be obtained without any additional constraints, which is consistent with the BARF and L2G works. In these methods, no additional loss functions were introduced apart from those used for NeRF training.
For the second case, when the intrinsics are globally unique, the number of unknown variables is remaining five. However, $u_{0}$ and $v_{0}$ can be shared across multiple views, which means that accurate parameters can be obtained by adding only a few constraints. Assuming that we have $m$ viewpoints, and there are a total of $3m$ constraint equations, while the number of unknown parameters is $23m+$ 2. For example, in SC-NeRF, the researchers introduced the Projected Ray Distance (PRD) loss function to achieve joint optimization. NeRF--- - provides no additional constraints, but the authors emphasized the limitation of their method, which is only effective for forward-facing scenes. We believe this effectiveness benefits from the neural network’s strong approximation capabilities, but this ability cannot generalize well in the absence of sufficient constraints.
For the third case, there is almost no difference between multi-view and single-view. Each viewpoint has its own intrinsics and extrinsics. Assuming we have $m$ viewpoints, there are a total of $3m$ constraint equations, while the number of unknown parameters is $5m$. This means that at least two additional constraint equations must be added for each viewpoint to obtain accurate parameters.
Fig.3 illustrates the comparison between joint optimization of extrinsics alone and the joint optimization of both intrinsics and extrinsics, highlighting the differences.
In summary, joint optimization in multi-camera systems is not merely a matter of turning camera parameters into learnable variables and combining them with NeRF for backpropagation training. Depending on the assumptions of the camera parameters, different numbers of constraint equations may be required to ensure the feasibility of the joint optimization results.
Figure 3: Coupling issue between intrinsic and extrinsic parameters. 1) The first row illustrates the joint optimization for extrinsics and NeRF. In methods such as BARF, L2G-NeRF, where the intrinsic parameters are known, mitigating the issue of camera parameters coupling. 2) The second row showcases the joint optimization of both intrinsics and extrinsic, as outlined in Eq.(5). This procedure effectively represents a transformation involving scaling, rotation, and translation. $u_{0}$ and $v_{0}$ from intrinsics, along with $t_{1}$, $t_{2}$ and $t_{3}$ from extrinsics, coexist within $(\tilde{T}-{R^{-1}}T)$ and cannot be disentangled.
3.2.3   Solution to the Coupling Issue
Based on Sec 3.2.2, a straightforward solution to address the coupling issue is imposing additional constraints on the coupled variables. We introduce auxiliary images based constraints to separate $u_{0}$, $v_{0}$ and $t_{1}$, $t_{2}$, $t_{3}$ from each other.
Given a set of 3D points $\left\{{C_{i}}\right\}$ in world coordinate system, and knowing their corresponding coordinates $\left\{{c_{i}}\right\}$ in pixel coordinate system. We define a homography $H$ with size $3\times 4$, which is also called camera matrix, the transformation can be defined as:
$\left\{{c_{i}}\right\}=H\left\{{C_{i}}\right\}\quad\text{ with }H=K\left[R\left|T\right.\right]$ (11)
With the condition of $i\geq 6$ and ${{\left\|H\right\|}_{F}}=1$, where $F$ is Frobenius norm, $H$ can be solved by SVD decomposition. When $H$ is obtained, $K$ and $R$ can be solved by the first three columns of $H$.
Eq.(11) implies that if we can provide a sufficient number ($i\geq 6$) of 3D-2D correspondences, additional constraints on the intrinsics can be introduced, enabling their estimation. It is important to note that the coordinate system of 3D points provided may differ from that of the target scene requiring reconstruction. In such cases, the estimated extrinsics become meaningless, but the intrinsics can still be useful and shared across different coordinate systems.
The correspondences provides additional constraints on $u_{0}$ and $v_{0}$ in the intrinsics, ensuring that Eq.(10) can be solved, resolving the coupling issue.
Figure 4: Degenerate case in intrinsic parameters estimation. When a single AprilTag is present in the calibration image, obtaining a valid solution is not feasible. At least two AprilTags ensures the acquisition of camera intrinsic parameters.
3.3   Degenerate Case and Auxiliary Image Scheme
In this section, we first explain the degenerate case and discuss how to prevent it (Sec 3.3.1). Based on this principle, we propose an efficient auxiliary image acquisition scheme (Sec 3.3.2). By using auxiliary images, we can simultaneously resolve parameter coupling and avoid degeneration during the joint optimization.
3.3.1   Degeneration in Coupling Solution
As discussed in Sec 3.2.3, providing more than six ($i\geq 6$) 3D-2D correspondences for each camera theoretically resolves the coupling issue. However, during practical system validation, we found that these correspondences can not always guarantee valid constraints. According to [ref52], when the 3D points in space are distributed on the same plane, Eq.(11) only yields a general solution rather than a unique one. This circumstance is referred to as degeneration.
The degeneration imposes rules on the provided correspondences, requiring that these matches follow the principle of non-coplanar distribution. In classical calibration methods, such as the chessboard method, all corners lie on the same plane, which can not meet our requirements. To obtain matches that satisfy these conditions, we designed an AprilTag cube and developed an auxiliary image acquisition scheme.
3.3.2   Auxiliary Images Capture Scheme
We use a multi-camera system to capture images containing the AprilTag cube, providing the aforementioned 3D-2D matches. The cube has each face spray-painted with a different AprilTag label from the 36H11 family. AprilTag can provide five calibration points per plane, including four corners and a central point. Detecting more than two AprilTag patterns in each image can provide at least 10 points, which fulfills the condition of being non-coplanar. Besides, the cube is easy to machine, with low cost but high machining accuracy. As for calibration points detection, AprilTag supports an open source algorithm, which is stable and easy to deploy.
Fig.5 illustrates the procedure for acquiring auxiliary images, and two sets of auxiliary images need to be captured. The first set, referred to as $Pack1$, is used to unify the world coordinate systems across all cameras and provide initial extrinsics. The second set, referred to as $Pack2$, introduces additional constraints to resolve the coupling issue and avoid degenerate case. Since the necessity of the second set has already been addressed, we here focus on discussing the role and importance of the first set of images.
Figure 5: Details of calibration data Acquisition. Firstly, the cube is captured by all cameras within the shared field of view, with its center defined as the origin of the world coordinate system. We define the set of images captured during this step as $Pack1$. Secondly, to ensure that each camera captures at least two AprilTags, providing non-coplanar correspondences, the operator needs to randomly move the cube in front of each camera. Once a camera detects more than two AprilTags, the image can be saved. Each camera only needs to capture one such image. The set of images captured in this phase is defined as $Pack2$.
According to Eq.(10), after decoupling the camera parameters using the auxiliary images from $Pack2$, our challenges shift from the third case in Sec 3.2.2 to the first one. At this point, we encounter the same challenges as those presented in BARF and L2G-NeRF. Although these methods can optimize imperfect extrinsics with NeRF, they are still susceptible to suboptimal solutions when the parameters are poorly initialized, as demonstrated in Sec 4.3. To avoid these suboptimal solutions that lead to optimization failure, we utilize more robust constraints provided by $Pack1$, ensuring that our method is not compromised by initialization.
Additionally, our image acquisition scheme is highly adaptable to multi-camera systems that undergo frequent changes. If the position or orientation of any camera changes, we simply need to place the cube at the center of the area and recapture $Pack1$, while $Pack2$ can reuse previously captured results. When adding a new camera or replacing an existing one, a single image containing at least two AprilTags must be captured for each new camera, followed by the recapture of $Pack1$. We believe this manual effort is significantly less than that required for performing a complete system calibration and does not necessitate evaluating the quality of any calibration results.
3.4   Parameters Design and Training Details
In this chapter, we first introduce how the camera parameters are transformed into learnable variables and outline the loss functions used in our method (Sec 3.4.1). Next, we define the world coordinate system and explain its relationship with the two sets of auxiliary images (Sec 3.4.2). Finally, given the different convergence rates of camera parameter estimation and NeRF training, we propose a training sequential diagram to ensure the efficiency of the training process (Sec 3.4.3).
3.4.1   Parametrization and Loss Function
Figure 6: Training sequence diagram: Our training process consists of three stages: 1) Camera Parameter Initialization Stage: The goal in this stage is to obtain coarse intrinsic and extrinsic parameters. Rendering training is not conducted in this stage. 2) Global Optimization Stage: Due to the coupling issue during joint optimization, we retain the Intrinsic Reprojection branch as a decoupling constraint. We also employ the progressive alignment method proposed in BARF to achieve extrinsic parameters with higher accuracy than those obtained in the first stage. 3) Fine-tuning Stage: In this stage, continuing to optimize extrinsic parameters without the progressive alignment constraint can lead to divergence. We freeze the extrinsic parameters and only adjust the intrinsic parameters.
Assume that the multi-camera system contains $m$ cameras with unknown parameters, and ${{K}_{j}}$$(j=0,1,2,...,m)$ denotes the intrinsic matrix of each camera. We decompose ${{K}_{j}}$ into the initialization ${{K}_{j0}}$ and the adjustable weights $\Delta{{K}_{j0}}$, and also assume the axes of the image coordinate system are all perpendicular to each other:
We define the adjustable extrinsic parameters in the $se(3)$ space and then convert back to the $SE(3)$ space. Let $\left({{\alpha}_{j0}},{{\alpha}_{j1}},{{\alpha}_{j2}},{{\alpha}_{j3}},{{\alpha}_{j4}},{{\alpha}_{j5}}\right)$ be 6 adjustable parameters in the $se(3)$ and the corresponding matrix in the $SE(3)$ is denoted as $\left[{{R}_{j}}\left|{{T}_{j}}\right.\right]$ with size $3\times 4$. For every camera in the system, we rewrite Eq.(11) as:
$pd\_{{c}_{ij}}={{K}_{j}}\left[{{R}_{j}}\left|{{T}_{j}}\right.\right]{{C}_{ij}}$ (13)
where ${{C}_{ij}}$ is the 3D point provided by $Pack2$, $pd\_{{c}_{ij}}$ is the predicted pixel coordinate. We define the loss function for the intrinsics constrains as:
$los{{s}_{intr}}=\sum\limits_{i=0,j=0}^{i=n,j=m}{\left\|\frac{gt\_{{c}_{ij}}}{\sqrt{{{h}^{2}}+{{w}^{2}}}}-\frac{pd\_{{c}_{ij}}}{\sqrt{{{h}^{2}}+{{w}^{2}}}}\right\|_{2}^{2}}$ (14)
where $h,w$ represent the height and width of the image, and $gt\_{{c}_{ij}}$ denotes the ground-truth coordinate of ${c}_{ij}$, which can be detected in the images including the AprilTag cube.
Solving for the extrinsics with the condition of providing intrinsic constraints is a typical Perspective-n-Point (PnP) problem. As a typical method to solve PnP, Bundle Adjustment (BA) can be described as to optimize extrinsics minimizing the difference between the projected 3D scene points in the images and their corresponding 2D image feature points. The loss function in BA is defined as:
where $h,w$ also represent the height and width of the image, and $gt\_{{p}_{ij}}$ denotes the AprilTag corners detected in $Pack1$. $pd\_{{p}_{ij}}$ represents the reprojected coordinates of the AprilTag corners during the PnP optimization.
In addition to the loss terms for camera parameters, the last loss function is the photometric loss commonly used in NeRF, defined as:
$los{s_{rgb}}=\sum{\left\|{\hat{C}(r)-C(r)}\right\|_{2}^{2}}$ (16)
where $\hat{C}(r)$ represents the ground truth RGB colors for ray $r$, and $C(r)$ represents the RGB colors predicted by the NeRF method.
3.4.2   World coordinate system
Although Eq.(14) and Eq.(15) differ only in their coefficients, their implications are entirely different. Eq.(14) uses the auxiliary images from $Pack2$ to constrain the intrinsics and address the coupling issue, with the extrinsics’ world coordinate system centered at each AprilTag cube. Since the cube in $Pack2$ is moved randomly, its position is unrelated to the multi-camera system’s world coordinate system. Consequently, the extrinsics optimized from Eq.(14) are redundant parameters, corresponding to the $Extrinsics$ $Local$ in Fig.2, and are not involved in the NeRF training stage.
Eq.(15) utilizes the auxiliary images from $Pack1$ to constrain the extrinsics of all cameras and to serve as the initial setup for these parameters. We define the coordinate system in $Pack1$ as the world coordinate system for the multi-camera system. In this scenario, the optimized extrinsics represent the extrinsics of each camera within the system and are integrated into the joint optimization with NeRF.
The origin of the world coordinate system is defined at the geometric center of the cube. The X-axis points towards the AprilTag labeled with number 1, the Y-axis points towards number 2, and the Z-axis points towards number 4.
Figure 7: Multi-camera image acquisition system in real-world and captured images containing AprilTags. The left side details the dimensions of our experimental site and the camera setup, utilizing a total of 88 cameras across the scene. The right side presents the real-world scene data we collected, divided into two sets. Consistent with the descriptions in Fig.2, $Pack1$ represents the Global package, while $Pack2$ represents the Local package. Because the site covered by the cameras is also the scene for NeRF representation, we use the images in $Pack1$ for training.
3.4.3   NeRF with Bundle Adjustment
Lin et al. first proposed Bundle-Adjusting Neural Radiance Fields (BARF), which addresses the problem of achieving NeRF training with imperfect camera extrinsics. One of the key advantages of BARF is its independence from requiring extra information. It only needs simple adjustments to the coarse-to-fine NeRF architecture to achieve the joint optimization for extrinsics and NeRF.
We attempted to extend this method by incorporating the estimation of intrinsics into the BARF framework. As a downstream component in the workflow, the rendering network is susceptible to fluctuations in upstream inputs, suggesting that training the rendering network could be ineffective when camera parameters change drastically. Moreover, the number of optimized parameters in the photometric loss branch far exceeds that of the intrinsic and extrinsic reprojection loss branches, which prolongs the training duration. Therefore, initializing camera parameters and subsequently refining them during rendering network training emerges as an efficient approach.
The training sequence is illustrated in Fig.6. During the camera parameters initialization stage, coarse camera parameters are obtained using the calibration data $Pack1$ and $Pack2$. In the global optimization stage, we employ the intrinsic reprojection loss branch to impose constraints on the principal point coordinates of the camera. We leverage the progressive alignment method from BARF to optimize the extrinsic parameters of the multi-camera acquisition system. In the fine-tuning stage, we freeze the extrinsic parameters and only refine the intrinsic parameters to obtain the final results.
3.5   Multi-Camera System Design and Dataset
In this section, we first introduce the multi-camera image acquisition system we constructed in the real world (Sec 3.5.1). Next, we present the proposed datasets, which includes both synthesis and real-world scenarios (Sec 3.5.2).
3.5.1   Multi-Camera Image Acquisition System
The motivation for MC-NeRF comes from the multi-camera image acquisition system constructed by ourselves. The dimensions of our site are a rectangular area measuring 9.0m $\times$ 6.0m $\times$ 2.4m. There are a total of 88 different cameras, each with independent camera parameters. Based on our experience, the calibration parameters of one camera cannot be shared with others, even though these cameras are claimed to be of the same model. The cameras are mounted on the truss in a uniform distribution, aligning with the setup commonly found in multi-camera image acquisition systems, such as facial imaging and motion capture systems. Fig.7. shows the detailed information of our site, as well as the scene data captured using our system.
3.5.2   Datasets of Multi-Camera Systems
Existing datasets, such as Synthesis, LLFF, and Mip-NeRF 360, typically capture scenes with a single camera, varying only in the extrinsics, which can not reflect the characteristics of multi-camera image acquisition systems. While some datasets for human reconstruction, such as HumanRF [ref63] and DNA-Rendering [ref64], use multi-camera systems, they fail to provide the auxiliary images. Our goal is to lower the barriers to using the NeRF method, and most current datasets are primarily designed to enhance rendering quality, which has motivated us to develop our own dataset.
Figure 8: Dataset Composition and Structure: The proposed datasets are divided into two parts: the left side illustrates the content and structure of the real-world dataset, while the right side displays the corresponding content and structure of the synthetic dataset. The real-world dataset comprises three scenes, each containing four components: the images of the scene to be reconstructed, two sets of auxiliary images, and calibration files for all cameras. These calibration files are critical for evaluating the accuracy of the estimated camera parameters. The synthetic dataset consists of four unique styles, each designed to simulate a different multi-camera acquisition system, with eight objects per style. Unlike the real-world dataset, the synthetic dataset additionally includes dedicated validation and test data.
Our public datasets consist of two parts: a synthetic dataset and a real-world scene dataset. For the synthetic dataset, we designed four packages based on commonly used multi-camera image acquisition systems. The camera distributions in these datasets include half-ball style (as used by the original NeRF), ball style, room style, and array style. The datasets contain a total of 32 groups, with 8 different objects for each style. The different sampling styles involve varying numbers of cameras, resulting in different numbers of training images.
The synthetic datasets are generated using Blender 3.3.3, and the intrinsic parameters for each camera within each style are different. For example, in the HalfBall Style acquisition system, there are a total of 100 cameras with varying intrinsic parameters. The training data is collected using these cameras according to their depicted positions. Similarly, the validation data is also collected using these cameras, but each camera’s position differs from that of the training set. For each style, the number of validation images is equal to the number of training images. The resolution of each image is $800\times 800$.
For the real-world scene dataset, we collect three sets of images using our custom-built system. Each scene includes 88 cameras with different parameters, and calibrated camera parameters are provided as ground truth for each camera. The resolution of each image is $960\times 540$. Both synthetic and real-world datasets include two sets of auxiliary images to validate our method. The details of the datasets are illustrated on Fig.8
It’s worth noting that due to the images being collected from cameras with different intrinsic parameters, determining which camera to use for generating the test data is challenging. To showcase the performance of the proposed method across various camera intrinsic parameters, we generate the 200 test images in each group by continuously varying the intrinsic parameters along the camera’s motion trajectory for synthetic dataset. Detailed sampling procedures and parameter variations have been provided on the project website.
4   EXPERIMENTS
All of our experiments were conducted on an NVIDIA RTX 3090 with 24GB of memory, using PyTorch version 1.12.1. For all comparative experiments, we utilized the publicly available source code from their respective papers. To facilitate the replication of our experiments, we did not adjust the hyperparameters of the open-source code, except for modifying the data loading workflow. The goal of the experiments is to answer the following questions:
  1. 1.
    How does the coupling issue affect the joint optimization process? In most methods that jointly optimize camera parameters and NeRF, 2D image alignment experiment is an effective way to validate the performance of joint optimization. How does the proposed method perform in the 2D image alignment experiments?
  2. 2.
    Sec 3.3.1 highlights that if the selected 3D points lie on the same plane, a degenerate case could occur. Does the proposed image acquisition scheme avoid this issue?
  3. 3.
    The author claims that $Pack1$ can be used to constrain the extrinsics and provide initial extrinsics. However, theoretically, $Pack2$ is already capable of providing constraints for solving all parameters. Is $Pack1$ necessary? Can the existing methods achieve accurate results if the extrinsics are initialized with random values?
  4. 4.
    How does the performance of the proposed method compare to other joint optimization approaches? Existing methods typically rely on straightforward combined strategies, such as using off-the-shelf SFM libraries to provide initial camera parameters. How does the performance of the proposed method compare to this strategy?
Figure 9: The performance of the 2D image alignment experiment. We visualize the optimized warps (top row), the patch reconstructions in corresponding colors (middle row), and recovered image representation from ${f_{mlp}}$ (bottom row). MC-NeRF is able to recover accurate alignment and high-fidelity image reconstruction. This indicates that joint optimization of intrinsics, extrinsics and NeRF involves more than merely converting these parameters into learnable variables and applying backpropagation.
4.1   Exp.1 2D Alignment for Coupling Issue
@@ 4   EXPERIMENTS
2D neural image alignment problem can be described as: let $x$ be the 2D pixel coordinates, we aim to optimize a 2D neural field parameterized as the weights of a multilayer perceptron (MLP) ${f_{mlp}}$:
@@
while also solving for geometric transformation as $T=[R{\rm{|}}T]$, and $[R{\rm{|}}T]$ is the rigid rotation and translation in 2D space.
@@
BARF 85.6157 14.2270 0.5149 0.6641
L2G-NeRF 55.3545 20.5440 0.7884 0.3364
MC-NeRF(w/o) 55.3096 15.9662 0.5923 0.6268
MC-NeRF(ours) 0.4084 40.8309 0.9828 0.0318
@@
$a_{1}$, $a_{2}$, $b_{1}$, $b_{2}$ are four parameters to be optimized, used to simulate inaccurate $f_{x}$, $f_{y}$, $u_{0}$, $v_{0}$ in intrinsics. It’s important to note that the term $T$ in Eq.(17) is defined as a homography matrix in original experiment. However, since the transformation requirements for 3D space in the NeRF series methods involve only rotation and translation, we redefine $T$ as $[R{\rm{|}}T]$ instead.
We choose the same image "cat" as BARF used from ImageNet [ref51]. We select five patches sampled from the original image, apply linear transformations and rigid perturbations, and optimize Eq.(18) to find the transformation $L$ and $T$ for each patch. Simultaneously, we train the neural network ${f_{mlp}}$ to represent the entire image. In all test models, a 6-layer fully connected network with 256 neurons each is respectively defined. The training epochs are set to 50,000 to ensure convergence for all methods. The training process is demonstrated on our project webpage.
Figure 10: Comparison results of estimated intrinsics with different numbers of AprilTags. With at least two AprilTags included in the auxiliary image, we can achieve the expected results. However, when including only one AprilTag, degeneration case is showed up, leading to poor results. Specifically, the black curve in $E1$ fails to converge to zero, while the blue and red curves both converge near zero. This means that when the selected 3D points are distributed on the same plane, degeneration occur, which is consistent with our conclusion.
The joint optimization of additional intrinsics can be equivalent to adding an additional linear transformation $L$ to the original rigid transformation. To decouple this linear transformation from rigid transformation, we randomly selected six pairs of calibration points as constraints for each patch. This constraint is analogous to the correspondences obtained by AprilTag cube in 3D space.
The render results are shown in Fig.9. The inclusion of linear transformations leads to both BARF and L2G-NeRF being unable to produce accurate results, which indicates that the coupling issue indeed exists. Even in cases where matches are not provided, proposed method also faces challenges in achieving precision results. With the condition of using calibration points for constraints, the accurately estimation of $L$ and $T$ can be obtained.
TABLE 2: Quantitative results of Exp.2. The loss items on the left represents the errors when estimating intrinsics using images containing different numbers of AprilTags. It can be observed that when using only one AprilTag for estimation (Dataset A), accurate intrinsics cannot be obtained, leading to a degenerate case.
@@
The quantitative results are shown in Table 1. $CornerError$ measures the absolute difference between the predicted and ground truth coordinates. This metric is calculated for the four vertices of a patch after applying the optimized $L$ and $T$. PSNR, SSIM, and LPIPS represent the rendering quality metrics for each patch. It can be seen that MC-NeRF achieved the best performance after decoupling constraints were applied.
4.2   Exp.2 Capture Scheme for Degeneration Case